OpenStudy (anonymous):

Write an equation for the line in point-slope and slope-intercept forms. Passes through (1, 2) and has a slope of 7. Can someone please show me how to do this?

OpenStudy (anonymous):

Do you know what point slope form is?\[y-ycoord=slope(x-xcoord)\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Passes through (1, 2) and has a slope of 7. That's the question and I have three more like it, can you please show me how to solve it?

OpenStudy (anonymous):

smeone please help?

OpenStudy (anonymous):

Plug in the numbers into my point slope equation

OpenStudy (anonymous):

wait I'm confused?

OpenStudy (anonymous):

can you show me how?

OpenStudy (texaschic101):

point slope form is : y - y1 = m(x - x1).... set 1 (1,2)--with x1 = 1 and y1 =2 slope (m) = 7 sub your info into the formula... y - y1 = m(x - x1) y - 2 = 7(x - 1) <--this is point slope form Now you need slop intercept form... slope intercept form is : y = mx + b your points -- (1,2)....your slope (m) = 7 now plug in your values... y = mx + b 2 = 7(1) + b 2 = 7 + b 2 - 7 = b -5 = b so your equation is : y = 7x - 5 <--slope intercept form

OpenStudy (anonymous):

Thank you so much I think I got the hang of it!!!!!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

thats good.....it can be confusing :)